# CalibRefine: Deep Learning-Based Online Automatic Targetless LiDAR–Camera Calibration with Iterative and Attention-Driven Post-Refinement

### v5

@@ I Introduction
Given these trade-offs, a fully automatic, targetless, and online calibration paradigm combines the most desirable attributes—removing cumbersome calibration objects, eliminating the need for human intervention, and adapting in real time to changing environments. However, despite its appeal, such a paradigm remains highly challenging, and existing approaches in this category exhibit various limitations. Motion-based methods often rely on additional hardware or specific sensor motion constraints, limiting real-world applicability [26]. Hand-eye calibration, while classical, demands multiple accurate sensor poses, rendering it unsuitable for static or dynamically constrained setups [27]. Edge-based methods suffer from unreliable edge matching across modalities, since object boundaries differ significantly across LiDAR and camera data [28, 29, 16]. Mutual information-based methods are similarly unreliable due to varying LiDAR reflectance and camera illumination sensitivities [30, 31]. Recent deep learning–based methods directly regress calibration parameters (e.g., RegNet [23] and its variants [24, 2, 17, 32, 33, 34]) but require initial manual calibrations or projected depth maps, and suffer from limited generalization and computational overhead, making them less suitable for real-time applications [17].
@@
To address these challenges, we propose CalibRefine, a fully automatic, targetless, online LiDAR–camera calibration framework. Our method directly processes raw LiDAR point clouds and camera images without initial calibration matrices or complex preprocessing. First, we leverage established object detection algorithms—YOLOv8 for camera data and an octree-based DBSCAN approach for LiDAR—to identify individual objects. A novel Common Feature Discriminator then matches these cross-sensor object instances by learning relative positions, appearance embeddings (using ResNet [35] for camera images and PointNet++ [36] for LiDAR data), and semantic class information, forming reliable cross-sensor correspondences. Recognizing potential inaccuracies from initial matching, we further enhance calibration through two online refinement stages. The iterative refinement incrementally optimizes calibration using accumulated correspondence data, while the attention-based refinement employs a Vision Transformer [37] and cross-attention mechanisms to correct for non-planar distortions and depth variations, further improving calibration accuracy. Crucially, our framework bypasses the pitfalls of direct matrix regression and the need for projected LiDAR maps, and eliminates the reliance on heuristic preprocessing or manually labeled calibration matrices, offering a more principled, data-driven pipeline that is both computationally efficient and adaptable in real-time. By integrating domain-specific mature object detection methods, a reliable discriminator to identify cross-sensor correspondences, and dual-stage refinement, our approach bridges the existing research gap, achieving a stable and accurate LiDAR–camera calibration that is truly automatic, targetless, and online. Our contributions are summarized as follows:
@@ III-A Problem Formulation
where ($\hat{P_{l}}=(\hat{u},\hat{v})$) is the projection of a point $P_{l}=(x,y)$ in the LCS onto the camera image plane PCS. Notably, the objects or points on the 2D LiDAR plane and those on the camera image plane are derived from objects or points lying on a common plane (e.g., the ground plane), as shown in Fig. 4. This alignment justifies the use of 2D Homography for LiDAR–camera calibration, as it can be considered a planar homography induced by the common plane [45]. To solve for the Homography matrix, a set of $N$ points in the LCS and their corresponding points in the PCS is required. Although 4 points are theoretically sufficient, using more points allows optimization of the solution via a cost function that minimizes the geometric reprojection error [46, 44]. This error, which measures the alignment between $N$ projected LiDAR points and their corresponding 2D image pixel points ($P_{p}=(u,v)$), can be quantified either as the Average Euclidean Distance (AED) or as the Root Mean Square Error (RMSE):
$\mathcal{E}_{\text{AED}}=\frac{1}{N}\sum_{i=1}^{N}\left\|P_{p}^{i}-\hat{P}_{l}^{i}\right\|_{2}=\frac{1}{N}\sum_{i=1}^{N}\sqrt{(u^{i}-\hat{u}^{i})^{2}+(v^{i}-\hat{v}^{i})^{2}}$ (2)
$\mathcal{E}_{\text{RMSE}}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left\|P_{p}^{i}-\hat{P}_{l}^{i}\right\|_{2}^{2}}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left[(u^{i}-\hat{u}^{i})^{2}+(v^{i}-\hat{v}^{i})^{2}\right]}$ (3)
@@ III-B Method Overview
We aim to develop a framework for LiDAR–camera online automatic targetless calibration that reduces human intervention, streamlines sensor integration, and ensures high precision in LiDAR–camera fusion applications. Our method comprises the following stages, as shown in Fig. 1: In Stage 1, established LiDAR and camera detectors are used to extract objects from each sensor’s data, capturing their center positions (bounding box centers for the camera detections and cluster centers for the LiDAR point cloud). These objects serve as training samples for a Common Feature Discriminator, which determines whether an image object and a LiDAR object correspond to the same entity by learning and comparing three distinct features (Relative Positions, Appearance Embeddings, and Classification Information) before concatenating them and passing them through a feed-forward neural network (FFN) classifier. In Stage 2, a homography transformation is applied to generate a coarse initial calibration matrix $H$ based on the identified object pairs, establishing a preliminary correspondence between the LiDAR and camera detections. Stage 3 refines this initial matrix iteratively by projecting LiDAR data onto the camera plane and creating additional point correspondences according to distance criteria, leading to a more precise calibration. Finally, in Stage 4, attention-based refinement employs a Vision Transformer to extract global distortion features from images, mitigating issues arising from non-planar surfaces, depth variations, and the absence of intrinsic matrix-based rectification. A cross-attention network computes weighted interactions among image pixels (queries), projected LiDAR points (keys), and LiDAR points (values), capturing more accurate correspondences between the two data sources. The model then learns a correction matrix $H^{\Delta}$ to refine the initial calibration, yielding an improved matrix $H^{*}$ for superior LiDAR–camera alignment.
@@ III-D Homography-based Calibration Matrix Estimation
Algorithm 1 Iterative Refined Calibration
Input: $F=\{(L_{i},C_{i})\}$ for $i=1\ldots T$trianglerightLiDAR and camera points for $T$ frames
     $H_{0}$trianglerightInitial calib. matrix from CFD
     $n\in\mathbb{Z}$trianglerightFrames to accumulate before recalib.
     $Bsz$trianglerightBlock size for sampling
Output: $H^{*}$trianglerightRefined calibration matrix
1:$H_{best}\leftarrow H_{0}$trianglerightSet initial matrix as best
2:$A\leftarrow[\ ]$trianglerightList to store pairs
@@
4:     $P^{L}_{i}\leftarrow proj(pts(L_{i}),H_{best})$trianglerightProject LiDAR points onto the camera plane
5:     $m\leftarrow GBMatch(P^{L}_{i},pts(C_{i}))$trianglerightGreedy Match points
6:     $s\leftarrow BBSample(m,Bsz)$trianglerightBlock-based sampling
7:     $A.\text{append}(s)$trianglerightAccumulate new pairs
@@
9:         $H_{new}\leftarrow Recalib(A)$trianglerightRecalibrate using accumulated pairs
10:         $e_{old}\leftarrow RepErr(H_{best},A)$
11:         $e_{new}\leftarrow RepErr(H_{new},A)$
12:         if $e_{new}<e_{old}$ then
13:              $H_{best}\leftarrow H_{new}$trianglerightUpdate matrix if error reduces
@@
17:$H^{*}\leftarrow H_{best}$trianglerightFinal refined calibration matrix
@@ III-E Iterative Refinement Process
Each projected point $(\hat{u}_{j},\hat{v}_{j})$ is then compared with the camera-detected object centers in $C_{i}$. A greedy bipartite matching algorithm [50] is used to associate each projected LiDAR point with its nearest camera detection (if one exists) based on a distance measure $d\bigl{(}(\hat{u},\hat{v}),(u,v)\bigr{)}$. Let $\mathcal{M}_{i}$ be the associated candidate point pairs set from frame $i$. Only candidate point pairs that fall within unoccupied or sufficiently distinct grid regions—determined by our block-based sampling strategy (Fig. 3)—are retained to form a filtered set ~$\widetilde{\mathcal{M}}_{i}\subseteq\mathcal{M}_{i}$, which is then incorporated into the accumulated set $\mathcal{A}$ via
@@
TABLE I: Performance of Common Feature Discriminator
Metric Dataset 1 Dataset 2
Binary Classification Accuracy (%) 98.00 92.50
Image Classification Accuracy (%) 82.80 72.00
LiDAR Classification Accuracy (%) 87.34 85.50
@@ III-F Attention-based Refinement Process
Our approach leverages a Vision Transformer (ViT) to capture global distortion features. Given an image $\mathbf{I}$ partitioned into patches $\{\mathbf{p}_{1},\dots,\mathbf{p}_{m}\}$, each patch is encoded into a token $\mathbf{t}_{i}=f_{\mathrm{ViT}}(\mathbf{p}_{i})$. The ViT applies multi-head self-attention,
@@
with token dimension $d$, thereby aggregating global information to reveal distortion patterns that simple planar models cannot capture.
Simultaneously, a cross-attention mechanism establishes precise correspondences between LiDAR and camera data. Image points $\{(u_{i},v_{i})\}_{i=1}^{N}$ generate queries $\mathbf{Q_{c}}$, while LiDAR points projected by $\mathbf{H}$ yield keys $\mathbf{K_{c}}$ and the original 3D coordinates $\{(x_{j},y_{j},z_{j})\}_{j=1}^{M}$ serve as values $\mathbf{V_{c}}$. The cross-attention output is computed as
@@
effectively linking each image point with its corresponding LiDAR feature.
The global features from the ViT and the detailed correspondence information from cross-attention are concatenated to form a unified feature vector,
@@
which is then processed by additional layers $g(\cdot)$ to regress a 9-dimensional vector $\bm{\theta}$. This vector is reshaped to obtain the correction matrix,
@@
resulting in the final refined homography $\mathbf{H}^{*}=\mathbf{H}\times\mathbf{H}^{\Delta}$. During training, a self-supervised loss minimizes the reprojection error between $\mathbf{H}^{*}$-projected LiDAR points and their corresponding image coordinates, guiding $\mathbf{H}^{\Delta}$ to correct any residual misalignments.
@@ IV-C2 Fine Calibration with Iterative Refinement
Tables III and IV detail the reprojection error evolution (using the $\mathcal{E}_{\text{AED}}$ metric) at different frame intervals (with an interval of 100 frames in our implementation) for Datasets 1 and 2. In each interval, the algorithm determines whether the newly computed homography matrix (New) provides a tighter alignment than the previously best-known matrix (Best); if so, it updates the calibration accordingly. Fig. 8(a) and 8(b) visualize these updates, where the blue line denotes the error obtained from the newly recalibrated matrix in each iteration, and the orange line tracks the evolving best-known solution. Not every recalibration step yields an improvement—reflecting the inherent noise and variability of real-world data—but key frame intervals (e.g., 0–100 for Dataset 1 and 0–1000 for Dataset 2) demonstrate significant error reductions, confirming that the iterative approach converges toward a more accurate solution over time. These updates demonstrate the iterative optimization process’s ability to adaptively refine the calibration as additional data and correspondences become available, ultimately enabling the iterative refinement to achieve significantly higher accuracy compared to the initial coarse calibration (as shown in Fig. 10(b)).

### v6

@@ I Introduction
Given these trade-offs, an automatic, targetless, and online calibration paradigm combines the most desirable attributes—removing calibration objects, eliminating human intervention, and adapting to changing environments. However, despite its appeal, such a paradigm remains highly challenging, and existing approaches in this category exhibit various limitations. Motion-based methods often rely on additional hardware or specific sensor motion constraints, limiting real-world applicability [26]. Hand-eye calibration, while classical, demands multiple accurate sensor poses, rendering it unsuitable for static or dynamically constrained setups [27]. Edge-based methods suffer from unreliable edge matching across modalities, since object boundaries differ significantly across LiDAR and camera data [28, 29, 16]. Mutual information-based methods are similarly unreliable due to varying LiDAR reflectance and camera illumination sensitivities [30, 31]. Recent deep learning–based methods directly regress calibration parameters (e.g., RegNet [23] and its variants [24, 2, 17, 32, 33, 34]) but require initial manual calibrations or projected depth maps, and suffer from limited generalization and computational overhead, making them less suitable for real-time applications [17].
@@
To address these challenges, we propose CalibRefine, a fully automatic, targetless, online LiDAR–camera calibration framework, as shown in Fig. 1 and 2. Our method directly processes raw LiDAR point clouds and camera images without initial calibration matrices or complex preprocessing. First, we leverage established object detection algorithms—YOLOv8 for camera data and an octree-based DBSCAN approach for LiDAR—to identify individual objects. A novel Common Feature Discriminator (CFD) then matches these cross-sensor object instances by learning relative positions, appearance embeddings (using ResNet [35] for camera images and PointNet++ [36] for LiDAR data), and semantic class information, forming reliable cross-sensor correspondences. Recognizing potential inaccuracies from initial matching, we further enhance calibration through two online refinement stages. The iterative refinement incrementally optimizes calibration using accumulated correspondence data, while the attention-based refinement employs a Vision Transformer [37] and cross-attention mechanisms to correct for non-planar distortions and depth variations, further improving calibration accuracy. Crucially, our framework bypasses the pitfalls of direct matrix regression and the need for projected LiDAR maps, and eliminates the reliance on heuristic preprocessing or manually labeled calibration matrices, offering a more principled, data-driven pipeline that is both computationally efficient and adaptable in real-time. By integrating domain-specific mature object detection methods, a reliable discriminator to identify cross-sensor correspondences, and dual-stage refinement, our approach bridges the existing research gap, achieving a stable and accurate LiDAR–camera calibration that is truly automatic, targetless, and online. Our contributions are summarized as follows:
@@ III-A Problem Formulation
[…]
$\begin{split}\mathcal{E}_{\text{AED}}&=\frac{1}{N}\sum_{i=1}^{N}\left\|P_{p}^{i}-\hat{P}_{l}^{i}\right\|_{2}\\ &=\frac{1}{N}\sum_{i=1}^{N}\sqrt{(u^{i}-\hat{u}^{i})^{2}+(v^{i}-\hat{v}^{i})^{2}},\end{split}$ (2)
or as the Root Mean Square Error (RMSE):
$\begin{split}\mathcal{E}_{\text{RMSE}}&=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left\|P_{p}^{i}-\hat{P}_{l}^{i}\right\|_{2}^{2}}\\ &=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left[(u^{i}-\hat{u}^{i})^{2}+(v^{i}-\hat{u}^{i})^{2}\right]}.\end{split}$ (3)
@@ III-B Method Overview
We aim to develop a framework for LiDAR–camera online automatic targetless calibration that reduces human intervention, streamlines sensor integration, and ensures high precision in LiDAR–camera fusion applications. Our method comprises the following stages (see Fig. 1 and 2): In Stage 1, the established LiDAR and camera detectors are utilized to detect and extract objects from their respective sensor data. Each detected object is extracted, including its center position (the bounding box center for the camera detection and the cluster center for the LiDAR point cloud). These objects are then used to train a Common Feature Discriminator, which determines whether an image object and a LiDAR object correspond to the same entity. To achieve this, the discriminator learns and compares three distinct features: Relative Positions, Appearance Embeddings, and Classification Information. These features are concatenated and passed through a feed-forward neural network (FFN) classifier, which outputs a decision on whether the objects from the two sensors are the same or different. Stage 2 involves solving the calibration matrix using the identified object pairs. A homography transformation is applied to generate a coarse initial calibration matrix ($H$), which establishes a preliminary correspondence between objects detected by the LiDAR and camera. In Stage 3, an iterative refinement-based fine calibration is performed to refine the calibration matrix, considering that the Common Feature Discriminator may not precisely match all corresponding objects. The initial calibration matrix from the previous stage is used to project LiDAR data onto the camera plane, enabling the selection and construction of additional point correspondences based on distance criteria. These newly established correspondences are then used to achieve more precise calibration through iterative refinement. Finally, in Stage 4, attention-based refinement employs a Vision Transformer to extract global distortion features from images, addressing the limitations of homography calibration caused by non-planar surfaces and depth variations. This also compensates for errors introduced by the absence of camera intrinsic matrix-based rectification. Furthermore, it integrates a cross-attention network to compute weighted interactions between image pixels (as queries), projected LiDAR points (as keys), and LiDAR points (as values), thereby capturing more accurate correspondences between LiDAR and camera data points. The model fundamentally learns and generates a correction matrix ($H^{\Delta}$) to refine the initial calibration, resulting in an improved matrix ($H^{*}$) for better LiDAR–camera alignment.
@@ III-D Homography-based Calibration Matrix Estimation
Algorithm 1 Iterative Refined LiDAR–Camera Calibration
[…]
@@
[…]
@@
[…]
12:   if $\mathcal{E}_{new}<\mathcal{E}_{old}$ then
[…]
@@
[…]
@@ III-E Iterative Refinement Process
Each projected point $(\hat{u}_{j},\hat{v}_{j})$ is then compared with the camera-detected object centers in $C_{i}$. A greedy bipartite matching algorithm [50] is used to associate each projected LiDAR point with its nearest camera detection (if one exists) based on a distance measure $d\bigl{(}(\hat{u},\hat{v}),(u,v)\bigr{)}$. Let $\mathcal{M}_{i}$ be the associated candidate point pairs set from frame $i$. Only candidate point pairs that fall within unoccupied or sufficiently distinct grid regions—determined by our block-based sampling strategy (Fig. 4)—are retained to form a filtered set $\widetilde{\mathcal{M}}_{i}\subseteq\mathcal{M}_{i}$, which is then incorporated into the accumulated set $\mathcal{A}$ via
@@ III-F Attention-based Refinement Process
Our approach leverages a Vision Transformer (ViT) to capture global distortion features. Given an image $\mathbf{I}$ partitioned into $m$ patches $\{\mathbf{p}_{1},\dots,\mathbf{p}_{m}\}$, each patch is encoded into a token $\mathbf{t}_{i}=f_{\mathrm{ViT}}(\mathbf{p}_{i})$. The ViT applies multi-head self-attention; for one head,
@@
where $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{m\times d}$ are the query, key, and value matrices obtained by learned linear projections of $\{\mathbf{t}_{i}\}$, and $d$ is the token feature dimension. This aggregates global information to reveal distortion patterns that planar models cannot capture. A global ViT representation is then obtained by average pooling,
$\mathbf{z}_{\mathrm{ViT}}\;=\;\frac{1}{m}\sum_{i=1}^{m}\mathbf{t}_{i}\in\mathbb{R}^{d}.$
Simultaneously, a cross-attention mechanism establishes precise correspondences between LiDAR and camera data. Image points $\{(u_{i},v_{i})\}_{i=1}^{N}$ produce queries $\mathbf{Q_{c}}\in\mathbb{R}^{N\times d}$; LiDAR points, projected to the image by the calibrated LiDAR-to-image matrix $\mathbf{H}$, yield keys $\mathbf{K_{c}}\in\mathbb{R}^{M\times d}$, and their 3D coordinates $\{(x_{j},y_{j},z_{j})\}_{j=1}^{M}$ provide values $\mathbf{V_{c}}\in\mathbb{R}^{M\times d}$. The cross-attention output is
@@
which links each image point to its corresponding LiDAR point.
The global ViT representation $\mathbf{z}_{\mathrm{ViT}}\in\mathbb{R}^{d}$ and the cross-attention output $\mathbf{A}_{\mathrm{cross}}\in\mathbb{R}^{N\times d}$ are concatenated to form a unified feature vector,
@@
which is then processed by a regression head $g:\mathbb{R}^{2d}\!\to\!\mathbb{R}^{9}$ to regress a 9-dimensional vector $\bm{\theta}$. This vector is reshaped to obtain the correction matrix,
@@
resulting in the final refined homography $\mathbf{H}^{*}=\mathbf{H}\times\mathbf{H}^{\Delta}$. During training, a self-supervised reprojection loss minimizes the discrepancy between $\mathbf{H}^{*}$-projected LiDAR points and their matched image coordinates. Let $\tilde{\mathbf{u}}_{j}=\Pi(\mathbf{H}^{*};x_{j},y_{j},1)\in\mathbb{R}^{2}$ denote the homogeneous projection (with normalization) of the $j$-th LiDAR point and $\mathbf{u}_{\pi(j)}=(u_{\pi(j)},v_{\pi(j)})$ its matched image point; we optimize
$\mathcal{L}_{\mathrm{reproj}}\;=\;\frac{1}{M}\sum_{j=1}^{M}\bigl{\|}\tilde{\mathbf{u}}_{j}-\mathbf{u}_{\pi(j)}\bigr{\|}_{2}^{2},$
which guides $\mathbf{H}^{\Delta}$ to correct residual misalignment.
@@ IV-C1 Coarse Calibration with Common Feature
TABLE I: Performance of Common Feature Discriminator
Metric Dataset 1 Dataset 2
Binary Classification Accuracy (%) 98.00 92.50
Image Classification Accuracy (%) 82.80 72.00
LiDAR Classification Accuracy (%) 87.34 85.50
@@ IV-C2 Fine Calibration with Iterative Refinement
Tables III and IV detail the reprojection error evolution (using the $\mathcal{E}_{\text{AED}}$ metric) at different frame intervals (with an interval of 100 frames in our implementation) for Datasets 1 and 2. In each interval, the algorithm determines whether the newly computed homography matrix (New) provides a tighter alignment than the previously best-known matrix (Best); if so, it updates the calibration accordingly. Fig. 10(a) and 10(b) visualize these updates, where the blue line denotes the error obtained from the newly recalibrated matrix in each iteration, and the orange line tracks the evolving best-known solution. Not every recalibration step yields an improvement—reflecting the inherent noise and variability of real-world data—but key frame intervals (e.g., 0–100 for Dataset 1 and 0–1000 for Dataset 2) demonstrate significant error reductions, confirming that the iterative approach converges toward a more accurate solution over time. These updates demonstrate the iterative optimization process’s ability to adaptively refine the calibration as additional data and correspondences become available, ultimately enabling the iterative refinement to achieve significantly higher accuracy compared to the initial coarse calibration (as shown in Fig. 11(b)).
@@ IV-C3 Fine Calibration with Attention-based Refinement
TABLE VII: Ablation of CalibRefine Components on Two Datasets (✓: enabled, ✗: disabled)
Settings Modules Dataset 1 Dataset 2 CFD RANSAC BBSample GBMatch Iterative ViT Cross-Attn $\mathcal{E}_{\text{AED}}\!\downarrow$ $\mathcal{E}_{\text{RMSE}}\!\downarrow$ $\mathcal{E}_{\text{AED}}\!\downarrow$ $\mathcal{E}_{\text{RMSE}}\!\downarrow$ Coarse 134.74 114.64 36.39 28.72 + Iterative 95.89 74.10 28.35 23.09 + Attention 93.27 72.68 26.40 22.25
TABLE VIII: Computational Demand Analysis (per 1080p frame). GPU usage measured on an Nvidia 32GB V100S.
Stage Operation Latency (ms) GPU (%) Notes CFD inference Image det. + LiDAR clustering 153.0 74 Frame-wise Projection & matching LiDAR $\to$ image, GBMatch, BBSample 18.1 18 Scales with #points Iterative refinement Recalibration on accumulated pairs 445.0/N=4.45 12 Periodic; N=100N{=}100 Attention refinement ViT + Cross-Attn $\to H^{\Delta}$ 166.0 82 Frame-wise Overall FPS: $\frac{1000}{339.5}\approx\mathbf{2.95}$ GPU: 82% End-to-End
@@
IV-C4 Ablation Study and Computation Evaluation
[…]
Table VIII reports the per-frame computational profile at 1080p on an Nvidia 32 GB V100S (batch size 1). The end-to-end throughput is $\approx 2.95$ FPS with peak GPU utilization $\approx 82\%$. The runtime is dominated by Attention refinement (166.0 ms) and CFD inference (153.0 ms); Projection & matching is modest (18.1 ms) but grows with the number of pairs, and the amortized cost of Iterative refinement is small (about $4.45$ ms/frame with N=100N{=}100). In practice, scheduling the attention module periodically (or on detected distributional shifts) can improve effective FPS with limited accuracy loss, given the narrow gap between “+Iterative” and “+Attention” in Table VII.
IV-C5 Limitations
CalibRefine relies on object-level correspondences to bootstrap and refine calibration; therefore, it assumes the presence of foreground objects within the overlapping fields of view. As such, scenes lacking objects (e.g., empty roads) are not applicable and were not included in our evaluation. Our experiments focus on planar urban intersections with fixed, pole-mounted sensors, implicitly leveraging ground coplanarity through a 2D homography. Formal error bounds under non-planar conditions (e.g., ramps or multi-level structures) are not provided at this stage and will be explored in future work. The experimental scope is further limited to sensor vendor-recommended, highest-resolution settings and default deployment viewpoints; we do not sweep sensor parameters (e.g., LiDAR line count, camera resolution) or alternative placements. Although the pipeline is fully automatic and requires no human intervention, its current throughput is constrained ($\approx 2.95$ FPS). We anticipate that real-time performance can be improved through stronger hardware, an optimized C++/TensorRT implementation, or the lighter backbones.